if the perimiter of the rectangle is 60 and the width is 12 what is the length?
Answer by mananth(16946)

let the length be l
width is 12
perimeter = 2*(l+w)
60=2*(l+12)
divide by 2
60/2 = 2*(l+12) / 2
30=l+12
add -12
30-12=l+12-12
18=l the length.
..
Check
18+18+12+12=60